OpenStudy (anonymous):

An elephant’s heart puts out an average power of 2.5 W but like a human heart, it is only 10% efficient. How many Calories does the elephant need to ingest daily just to keep its heart pumping? (4190 J = 1 Cal).

OpenStudy (irishboy123):

start with daily **energy** output [at 10% efficiency]: \(2.5 W \times 24 \times 60 \times 60 = ???? \; J\)

OpenStudy (anonymous):

oh i see! thanks alot :D

OpenStudy (anonymous):

would u use the 10% in the calculations what so ever?

OpenStudy (irishboy123):

well, those numbers are only 10% of what it is really producing. so you need to do the sums and then gross up.

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

ok sorry once u find the joules u convert to calories, but u get such a small value? i think im doing something wrong :/